CC12:

PLEASE HELP ASAP IM STUCK ON THIS QUESTION Use the quadratic formula to determine the exact solutions to the equation. 2x^2−5x+1=0 Enter your answers in the boxes. x = or x =

CC12:

@Ultrilliam

MoonShadow:

The quadratic formula is x = (-b±√(b2-4ac) ) / (2a) This comes from completing the square on ax2+bx+c = 0. x = (5±√(25-4*2*1)) / (2*2) x = (5±√17) / 4 Your two solutions are x = (5 - √17) / 4 and x = (5 + √17) / 4

CC12:

Thank you so much

MoonShadow:

Np